OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y = 4x - 11 y = 4x + 11 y = 4x + 7 y = 4x - 7

OpenStudy (anonymous):

The answer is B. (y+1) = 4(x+3) y=4x+11. Just plug in

OpenStudy (anonymous):

@natalie135 can you help explain this to me?

OpenStudy (anonymous):

Ya so the equation for a line is: \[(y-y _{1})=m(x-x _{1})\] We plug in the points that were given: (-3, -1) for x1 and y1. The slope is 4, which corresponds to m. \[(y-(-1)=4(x-(-3))\] Make sense?